# FUIA: Model Inversion Attack against Federated Unlearning

### v2

@@ V-B1 Step 1: Gradient Separation
In the gradient separation step, the objective is to extract the “clean” gradient information of the data on the target client from the global model parameters obtained through FL. Suppose that client makes the unlearning request, and after the FU process, the data on the client has been entirely unlearned, meaning that the global model no longer contains their gradient information. Therefore, unlike sample unlearning, this step does not require focusing on the gradient changes in the unlearning process but only needs to separate the gradient information of client from the FL training process. Specifically, by calculating the parameter differences for each round and weighting them, we can obtain the “clean” gradient information of client , denoted as , as shown in the following formula:
@@ V-B3 Step 3: Gradient Inversion
represents the gradient of the constructed virtual sample , is the total variation regularization term used to control the smoothness of the reconstructed image and reduce noise and artifacts, is a hyperparameter that balances the regularization term, and is used to balance the influence of and . Through this optimization process, the attacker can construct a virtual sample whose features closely resemble the data on the target client, successfully revealing their feature information.
@@ V-C FUIA against Class Unlearning
Existing studies have shown that the removal of large amounts of data has a significant impact on the model parameters [32]. The training process of deep learning models essentially optimizes the model parameters to improve performance. When data from a specific class is deleted, the model parameters, particularly those in the output layer associated with that class, will change. This is because, after deleting the data points, the model no longer receives training signals from these data, which causes the weights of the output nodes related to that class to update less frequently and may even diminish or converge to zero. After the unlearning process removes data from a specific class, the server, acting as the attacker, can infer the characteristics of that class by comparing the weight differences in the output layer before and after the unlearning.
@@
where is a hyperparameter that balances the weight and bias differences. Based on the discrimination score, we can determine the forgotten class. The class with the highest score is considered the unlearned class. The class ID for unlearning can be determined as:
Through this process, the attacker can accurately identify the forgotten class, thereby leaking privacy information about the training data.
@@ VI Experiments and Analysis
In this section, we systematically validate the effectiveness of FUIA through extensive experiments. First, we elaborate on the experimental settings. Subsequently, we design comparative experiments for different types of FU tasks to comprehensively evaluate the attack performance of FUIA. Besides, we conduct an in-depth analysis of the experimental results.
@@ VI-A1 Datasets
In the experiments, we employed pretraining in the FL training process for sample unlearning and client unlearning tasks. Specifically, we randomly partitioned the dataset into two disjoint subsets, with 80 used to train a pretrained model and the remaining 20 serving as private data for individual clients. After pretraining, further training was conducted on client-specific data. This approach is widely adopted in FL as it enhances model performance. By leveraging large-scale data to capture generalized features, the pretrained model provides a more stable initialization for client fine-tuning, thereby reducing training time and computational overhead while improving convergence speed and generalization capability [35]. In the class unlearning task of FL, we do not adopt a pretraining approach. This decision stems from the fact that data from the target class to be unlearned may exist in both the training data of the pretrained model and the private datasets of the clients. Since our dataset partitioning is random and does not enforce a specific class to be client-private, it contradicts the fundamental assumption of pretraining, which assumes that class data can be shared between training and testing datasets.
@@ VI-A3 FL settings
We configure 10 clients, and in each training round, 50 of the clients are randomly selected to participate. Each participating client conducts 3 local training rounds on their own data. In the standard experiment, we use the FedAvg aggregation method to combine the updates from all clients, which is widely applied in FL and FU due to its ability to effectively balance the training contributions from all clients.
@@ VI-A4 FU settings
To comprehensively analyze the privacy vulnerabilities of FU, we selected two representative unlearning methods for each of the three FU types to evaluate the effectiveness of FUIA. Retraining, as the most fundamental and widely used method, is applicable to all FU types. Additionally, for class unlearning, we selected FUCDP [11] for evaluation; for client unlearning, we adopted FUEraser [9] as a baseline method; and for sample unlearning, we employed UnrollingSGD [37] on the target client for unlearning, followed by normal FL aggregation.
@@ VI-A5 Metrics
In the FUIA for sample unlearning and client unlearning, we use MSE (Mean Squared Error) and PSNR (Peak Signal-to-Noise Ratio) to evaluate the effectiveness of image reconstruction [38]. Specifically, MSE reflects the difference between the original image and the reconstructed image, with a smaller difference indicating a higher similarity between the two, while PSNR measures the signal-to-noise ratio of the reconstructed image, with a higher ratio indicating better image quality and more effective attacks. In the FUIA for class unlearning, we use prediction accuracy to assess the attack effectiveness. Specifically, we evaluate whether the predicted class label matches the actual label of the forgotten class data.
@@ VI-A6 Comparison Method and implementation
We adapt the MUIA [26] and apply it within the FL environment for comparative analysis. Specifically, we utilize the difference between the global models of FL before and after unlearning as the gradient information of the forgetting data for MUIA. All experiments were performed on a high-performance computing system equipped with an NVIDIA 3090 GPU (24 GB of VRAM) and Intel Xeon processors.
@@ VI-B1 Effectiveness of FUIA against Sample Unlearning
To simplify the experiment and more conveniently demonstrate the attack effectiveness, we start by exploring a relatively simple scenario. Specifically, to reduce computational cost and improve experimental efficiency, we set the sample size on each client to 8 and randomly select 1 data point for unlearning. In Section VII, we conduct additional experimental analysis regarding the number of samples per client and the amount of forgotten data to explore how these factors influence the attack performance. The results are shown in Figure 5, where we randomly select the reconstruction results of two clients for display. The figure illustrates the reconstruction performance of FUIA and MUIA against two FU methods, Retraining and UnrollingSgd, on the CIFAR-10 and CIFAR-100 datasets. To quantitatively evaluate the reconstruction performance, we also analyzed the results using the MSE and PSNR.
@@ VI-B2 Effectiveness of FUIA against Client Unlearning
AS shown in Figure 6, FUIA successfully reconstructs the data of the target client for both Retraining and FUEraser, indicating that client unlearning-based FU methods can lead to privacy leakage. Both in terms of image detail recovery and performance in metrics such as MSE and PSNR, FUIA shows significant effectiveness. Similar to the results for sample unlearning, the reconstruction performance of FUIA is clearly superior to that of MUIA, particularly in terms of image clarity and detail recovery. However, FUIA still shows some performance gaps when applied to Retraining compared to approximate FU methods like FUEraser. In conclusion, FUIA is highly effective in client unlearning attacks, capable of reconstructing relatively clear images under the current setup, leading to significant privacy leakage in FU scenarios. Furthermore, it outperforms other methods.
@@ VI-B3 Effectiveness of FUIA against Class Unlearning
| 3 | 10 | |
| 4 | 90 |
In this experiment, to better illustrate the effectiveness of the attack, we primarily conduct experiments using the CIFAR-10 dataset. In each experiment, we randomly select data from one class out of ten classes for unlearning and used FUIA to analyze the differences between the models before and after unlearning. By calculating the list of discrimination scores, we infer the labels of the unlearned class data. Furthermore, in Section VII, we conduct additional experiments on unlearning multiple classes to verify the broad applicability and robustness of FUIA.
The experimental results are shown in Figure 7, where we randomly selected two experimental instances to demonstrate the attack effectiveness of FUIA under the Retraining and FUCDP unlearning methods. The figure presents the discrimination scores for each data class, with red and blue bars representing the Retraining and FUCDP methods, respectively. In Figure 7(a), when unlearning data from class 0, FUIA accurately identifies class 0 as having the highest discrimination score, regardless of the unlearning method used. Similarly, in Figure 7(b), when unlearning class 9, FUIA consistently reveals class 9 as having the highest discrimination score, demonstrating the robustness of the attack. Furthermore, we measured the probability of FUIA correctly inferring the forgotten class, as shown in Table I. When unlearning a single class, FUIA achieves a 100 success rate in identifying the forgotten class. These results indicate that FUIA can directly infer the labels of unlearned data, significantly exacerbating privacy leakage risks.
@@ VII Ablation Study
In this section, we first conduct additional ablation experiments on the key variables involved in the previous experiments to further evaluate their impact on the performance of FUIA. Additionally, we thoroughly analyze the influence of different aggregation methods in FL on the attack effectiveness of FUIA.
@@ VII-A2 The Number of Forgotten Data on Each Client
We further investigate the impact of the number of forgotten data samples per client on the effectiveness of FUIA attacks. Specifically, we use the CIFAR-10 dataset and select 2 and 4 data samples for unlearning on each client, followed by an attack using FUIA. The experimental results are shown in Figure 9, where we present the attack outcomes for one randomly selected client. When each client forgets 2 data samples, FUIA remains effective in reconstructing the images, preserving certain original data features and causing partial privacy leakage. However, compared to the scenario where only 1 data sample is forgotten, the reconstructed images exhibit reduced clarity, with less distinct details and noticeable parameter deviations. When the number of forgotten samples increases to 4 per client, the attack performance of FUIA deteriorates further. While some reconstructed images still retain certain features of the forgotten data, they become significantly more blurred, and in some cases, it becomes impossible to fully reconstruct all forgotten samples, with some feature information completely lost. We hypothesize that this phenomenon is primarily due to the increasing difference in model parameters before and after unlearning as the number of forgotten data samples grows. Notably, since FUIA relies on estimated rather than exact gradient information, larger parameter variations lead to more ambiguous target gradients, thereby weakening the attack performance. In conclusion, increasing the number of forgotten data samples per client significantly reduces the effectiveness of FUIA.
@@ VII-B Ablation Study for FUIA against Client Unlearning
We further explore the impact of the number of data per client on the effectiveness of the FUIA attack. Specifically, to better present the experimental results, we selected CIFAR-10 as the dataset and set the number of data points per client to 8, then randomly chose one client for unlearning and performed the FUIA attack. The experimental results, shown in Figure 10, present the outcome of a randomly selected trial. It is evident from the figure that the effectiveness of the FUIA attack significantly decrease. Only a partial reconstruction of the forgotten client data is achieved, and the reconstructed image is relatively blurred. In terms of parameters, the MSE significantly increase compared to the previous results, while the PSNR clearly decrease. Objectively, it is considerably more challenging to attack client unlearning compared to sample unlearning. Unlearning an entire client causes a larger impact on the model parameters, introducing more significant disturbances. Additionally, accurately collecting gradient information for the entire client is a difficult task, which is also reflected in the previous MUIA experimental results, where MUIA almost failed to reconstruct any meaningful information from the target client. Overall, as the number of forgotten clients increases substantially, the effectiveness of the FUIA is significantly weakened.
@@ VII-C Ablation Study for FUIA against Class Unlearning
To further analyze the impact of unlearning multiple classes on the effectiveness of FUIA attacks, we investigate scenarios where multiple classes are forgotten in a single unlearning process. Specifically, we randomly select 2 and 3 classes for unlearning and then perform attacks using FUIA. The experimental results are presented in Figure 11, showing two trials. In Figure 11(a), when classes 0 and 9 are selected for unlearning, the discrimination scores computed by FUIA indicate that these two classes have the highest scores, successfully identifying the forgotten class labels. Similarly, in Figure 11(b), when unlearning 3 classes, FUIA is still able to accurately infer the forgotten class labels. Furthermore, we conducted extensive experiments to evaluate the attack performance of FUIA in multi-class unlearning scenarios. As shown in Table I, when 4 classes are unlearned, FUIA exhibits minor errors in label inference and fails to fully recover all forgotten class labels. This may be due to the correlation between the forgotten and remaining classes, leading to a small probability that some remaining classes have slightly higher scores than the forgotten ones. Overall, FUIA remains effective in attacking multi-class unlearning, but its accuracy decreases as the number of forgotten classes increases.
@@ VII-D Aggregation Method
We conduct a comparative analysis of the attack performance of FUIA when using FedAvg and FedSGD as aggregation methods. Specifically, we apply FUIA to attack sample unlearning experiments using the Retraining method aggregated by FedAvg and FedSGD. In the experiments, the attacker (server) obtains local model parameter information via FedAvg, while it can directly access local model training gradient information via FedSGD. The experimental results, shown in Figure 12, clearly indicate that FUIA achieves significantly better attack performance when FedSGD is used as the aggregation method compared to FedAvg. Moreover, the MSE and PSNR values further confirm this observation. This phenomenon aligns with our previous analysis. Since FUIA relies on gradient information for attacks, FedSGD enables the attacker to directly and accurately obtain the target gradients, reducing estimation errors and achieving better attack performance compared to FedAvg.
@@ VIII Possible defenses
In this section, we discuss two possible defense methods against FUIA. Since FUIA is essentially based on GIA in FL, we draw inspiration from related research [18, 15] to design two general defense strategies. In our experiments, we use the CIFAR-10 dataset and focus on sample unlearning to demonstrate the defense effectiveness, as the FUIA mechanisms for sample unlearning and client unlearning both rely on GIA.

### v3

@@ V-B1 Step 1: Gradient Separation
In the gradient separation step, the objective is to extract the “clean” gradient information of the data on the target client from the global model parameters obtained through FL. Suppose that client makes the unlearning request, and after the FU process, the data on client has been entirely unlearned, meaning that the global model no longer contains their gradient information. Therefore, unlike sample unlearning, this step does not require focusing on the gradient changes in the unlearning process but only needs to separate the gradient information of client from the FL training process. Specifically, by calculating the parameter differences for each round and weighting them, we can obtain the “clean” gradient information of client , denoted as , as shown in the following formula:
@@ V-B3 Step 3: Gradient Inversion
represents the gradient of the constructed virtual sample , is the total variation regularization term used to control the smoothness of the reconstructed image and reduce noise and artifacts, and is a hyperparameter that balances the regularization term. To balance the influence of and , we set and analyze it in detail in section VII. Through this optimization process, the attacker can construct virtual samples whose features closely resemble the data on the target client, successfully revealing their feature information.
@@ V-C FUIA against Class Unlearning
Existing studies have shown that the removal of large amounts of data has a significant impact on the model parameters [31]. The training process of deep learning models essentially optimizes the model parameters to improve performance. When data from a specific class is deleted, the model parameters, particularly those in the output layer associated with that class, will change [33]. This is because, after deleting the data points, the model no longer receives training signals from these data, which causes the weights of the output nodes related to that class to update less frequently and may even diminish or converge to zero. After the unlearning process removes data from a specific class, the server, acting as the attacker, can infer the characteristics of that class by comparing the weight differences in the output layer before and after the unlearning.
@@
where we set to equally balance the contributions of the weight and bias differences. Based on the discrimination score, we can determine the forgotten class. The class with the highest score is considered the unlearned class. The class id for unlearning can be determined as:
In cases where classes have been unlearned, we sort the discriminative scores diff_score and select the top- scores to obtain the labels of forgotten classes. Through this process, the attacker can accurately identify the forgotten class, thereby leaking privacy information about the training data.
@@ VI-A1 Datasets
In the experiments, we employed pre-training in the FL training process for sample unlearning and client unlearning tasks. Specifically, we randomly partitioned the dataset into two disjoint subsets, with 80 used to train a pre-trained model and the remaining 20 serving as private data for individual clients. After pre-training, further training was conducted on client-specific data. This approach is widely adopted in FL as it enhances model performance. By leveraging large-scale data to capture generalized features, the pre-trained model provides a more stable initialization for client fine-tuning, thereby reducing training time and computational overhead while improving convergence speed and generalization capability [35]. In the class unlearning task, we do not adopt a pre-training approach. This decision stems from the fact that data from the target class to be unlearned may exist in both the training data of the pre-trained model and the private datasets of the clients. Since our dataset partitioning is random and does not enforce a specific class to be client-private, it contradicts the fundamental assumption of pre-training, which assumes that class data can be shared between training and testing datasets.
@@ VI-A3 FL settings
We configure 10 clients, with 50 randomly selected to participate in each training round. Each participating client performs 3 local training rounds on its own data. In the standard experiment, we use FedAvg as the aggregation method, which is widely applied in FL and FU.
@@ VI-A4 FU settings
To comprehensively analyze the privacy vulnerabilities of FU, we select two representative unlearning methods for each of the three FU scenarios to evaluate the effectiveness of FUIA. Retraining, as the most fundamental method, applies to all FU scenarios and represents exact FU methods. Additionally, we selected widely studied representatives of approximate FU for testing. For class unlearning, we chose FUCDP [11] for evaluation; for client unlearning, we use FUEraser [9] as a comparative method; and for sample unlearning, we perform unlearning on the target client using UnrollingSGD [37] and conduct global model aggregation through the standard FL process.
@@ VI-A5 Metrics
In the FUIA for sample unlearning and client unlearning, we use MSE (Mean Squared Error) and PSNR (Peak Signal-to-Noise Ratio) to evaluate the effectiveness of image reconstruction [38]. Specifically, MSE reflects the difference between the original image and the reconstructed image, with a smaller difference indicating a higher similarity between the two, while PSNR measures the signal-to-noise ratio of the reconstructed image, with a higher ratio indicating better image quality and more effective attacks. For class unlearning, we use prediction accuracy to assess the attack effectiveness. Specifically, we evaluate whether the predicted labels match the actual labels of the forgotten classes.
@@ VI-A6 Comparison Method and Implementation
We adopt the MUIA [25] and apply it within the FL environment for comparative analysis. Specifically, we utilize the difference between the global models of FL before and after unlearning as the gradient information of the forgetting data for MUIA. All experiments are performed on a computing system equipped with an NVIDIA 3090 GPU (24 GB of VRAM) and Intel Xeon processors.
@@ VI-B1 Effectiveness of FUIA against Sample Unlearning
To conveniently demonstrate the effectiveness of the attack, we start by exploring a relatively simple scenario. Specifically, to reduce computational cost and improve experimental efficiency, we set the sample size on each client to 8 and randomly select 1 data point for unlearning. In Section VII, we conduct additional experimental analysis regarding the number of samples per client and the number of forgotten data to explore how these factors influence the attack performance. The results are shown in Figure 5, where we randomly select the reconstruction results of two clients for display. The figure illustrates the reconstruction performance of FUIA and MUIA against two FU methods, Retraining and UnrollingSgd, on the CIFAR-10 and CIFAR-100 datasets. To quantitatively evaluate the reconstruction performance, we also analyzed the results using MSE and PSNR.
@@ VI-B2 Effectiveness of FUIA against Client Unlearning
AS shown in Figure 6, FUIA successfully reconstructs the data of the target client for both Retraining and FUEraser, indicating that client unlearning-based FU methods can lead to privacy leakage. Both in terms of image detail recovery and performance in metrics such as MSE and PSNR, FUIA shows significant effectiveness. Similar to the results for sample unlearning, the reconstruction performance of FUIA is clearly superior to that of MUIA, particularly in terms of image clarity and detail recovery. However, FUIA still shows some performance gaps when applied to Retraining compared to approximate FU methods like FUEraser. In conclusion, FUIA is highly effective in client unlearning attacks, capable of reconstructing relatively clear images under the current setup, leading to significant privacy leakage in FU scenarios. Furthermore, it outperforms MUIA.
@@ VI-B3 Effectiveness of FUIA against Class Unlearning
| 3 | 20 | |
| 4 | 80 |
To better illustrate the effectiveness of the attack, we primarily conduct experiments using the CIFAR-10 dataset. In each experiment, we randomly select data from one class out of ten classes for unlearning and use FUIA to analyze the differences between the models before and after unlearning. By calculating the list of discrimination scores, we infer the labels of the unlearned class data. Furthermore, in Section VII, we conduct additional experiments on unlearning multiple classes to verify the broad applicability and robustness of FUIA.
The experimental results are shown in Figure 7, where we randomly select two experimental instances to demonstrate the attack effectiveness of FUIA under the Retraining and FUCDP unlearning methods. The figure presents the discrimination scores for each data class, with red and blue bars representing the Retraining and FUCDP methods, respectively. In Figure 7(a), when unlearning data from class 0, FUIA accurately identifies class 0 as having the highest discrimination score, regardless of the unlearning method used. Similarly, in Figure 7(b), when unlearning class 9, FUIA consistently reveals class 9 as having the highest discrimination score, demonstrating the robustness of the attack. Furthermore, we measure the probability of FUIA correctly inferring the forgotten class, as shown in Table I. When unlearning a single class, FUIA achieves a 100 success rate in identifying the forgotten class. These results indicate that FUIA directly infers the labels of unlearned data, significantly exacerbating privacy leakage risks.
@@ VII-A2 The Number of Forgotten Data on Each Client
We further investigate the impact of the number of forgotten data samples per client on the effectiveness of FUIA attacks. Specifically, we use the CIFAR-10 dataset and select 2 and 4 data samples for unlearning on each client, followed by an attack using FUIA. The experimental results are shown in Figure 9, where we present the attack outcomes for one randomly selected client. When each client forgets 2 data samples, FUIA remains effective in reconstructing the images, preserving certain original data features and causing partial privacy leakage. However, compared to the scenario where only 1 data sample is forgotten, the reconstructed images exhibit reduced clarity, with less distinct details and noticeable parameter deviations. When the number of forgotten samples increases to 4 per client, the attack performance of FUIA deteriorates further. While some reconstructed images still retain certain features of the forgotten data, they become significantly more blurred, and in some cases, it becomes impossible to fully reconstruct all forgotten samples, with some feature information completely lost. We assume that this phenomenon is primarily due to the increasing difference in model parameters before and after unlearning as the number of forgotten data samples grows. Notably, since FUIA relies on estimated rather than exact gradient information, larger parameter variations lead to more ambiguous target gradients, thereby weakening the attack performance. In conclusion, increasing the number of forgotten data samples per client reduces the effectiveness of FUIA.
@@ VII-B Ablation Study for FUIA against Client Unlearning
VII-B1 The value of
As previously discussed, not only represents the gradient information of the forgotten client but also contains noise introduced from other clients during the FL process. Therefore, it is used as an auxiliary term to enhance the gradient features of the target client. To balance the influence of and , we set to suppress noise interference and improve reconstruction quality. To further analyze the effect of on FUIA’s reconstruction performance, we conduct a set of experiments using the CIFAR-10 dataset. We set the number of data points per client to 1 and randomly choose one client for unlearning. The FUIA is then applied using different values of . As shown in Figure 10, when exceeds 0.1, the quality of the reconstructed images decreases. The MSE increases, and the PSNR decreases, indicating that excessive use of introduces detrimental noise. This supports our hypothesis that should be treated only as auxiliary information. On the other hand, when is set too low, the reconstruction metrics also degrade, demonstrating that does contain useful gradient information related to the forgotten client and should not be ignored. Based on these observations, we adopt to effectively utilize as an auxiliary term.
VII-B2 The Number of data on Each Client
We further explore the impact of the number of data per client on the effectiveness of the FUIA attack. Specifically, to better present the experimental results, we selected CIFAR-10 as the dataset and set the number of data points per client to 8, then randomly chose one client for unlearning and performed the FUIA attack. The experimental results, shown in Figure 11, present the outcome of a randomly selected trial. It is evident from the figure that the effectiveness of the FUIA attack significantly decrease. Only a partial reconstruction of the forgotten client data is achieved, and the reconstructed image is relatively blurred. In terms of parameters, the MSE significantly increase compared to the previous results, while the PSNR clearly decrease. Objectively, it is considerably more challenging to attack client unlearning compared to sample unlearning. Unlearning an entire client causes a larger impact on the model parameters, introducing more significant disturbances. Additionally, accurately collecting gradient information for the entire client is a difficult task, which is also reflected in the previous MUIA experimental results, where MUIA almost failed to reconstruct any meaningful information from the target client. Overall, as the number of forgotten clients increases substantially, the effectiveness of the FUIA is weakened.
@@ VII-C Ablation Study for FUIA against Class Unlearning
To further analyze the impact of unlearning multiple classes on the effectiveness of FUIA attacks, we investigate scenarios where multiple classes are forgotten in a single unlearning process. Specifically, we randomly select 2 and 3 classes for unlearning and then perform attacks using FUIA. The experimental results are presented in Figure 12, showing two trials. In Figure 12(a), when classes 0 and 9 are selected for unlearning, the discrimination scores computed by FUIA indicate that these two classes have the highest scores, successfully identifying the forgotten class labels. Similarly, in Figure 12(b), when unlearning 3 classes, FUIA is still able to accurately infer the forgotten class labels. Furthermore, we conducted extensive experiments to evaluate the attack performance of FUIA in multi-class unlearning scenarios. As shown in Table I, when 4 classes are unlearned, there is a 20 chance that FUIA can not fully infer the labeling of the forgotten category. This may be due to the correlation between the forgotten and remaining classes, leading to a small probability that some remaining classes have slightly higher scores than the forgotten ones. Overall, FUIA remains effective in attacking multi-class unlearning, but its accuracy decreases as the number of forgotten classes increases.
@@ VII-D Aggregation Method
We conduct a comparative analysis of the attack performance of FUIA when using FedAvg and FedSGD as aggregation methods. Specifically, we apply FUIA to attack sample unlearning experiments using the Retraining method aggregated by FedAvg and FedSGD. In the experiments, the attacker obtains local model parameter information via FedAvg, while it can directly access local model training gradient information via FedSGD. The experimental results, shown in Figure 13, clearly indicate that FUIA achieves significantly better attack performance when FedSGD is used as the aggregation method compared to FedAvg. Moreover, the MSE and PSNR values further confirm this observation. This phenomenon aligns with our previous analysis. Since FUIA relies on gradient information for attacks, FedSGD enables the attacker to directly and accurately obtain the target gradients, reducing estimation errors and achieving better attack performance compared to FedAvg.
@@ VIII Possible defenses
In this section, we draw inspiration from related research [39, 15] to design two general defense strategies. In our experiments, we use the CIFAR-10 dataset and choose sample unlearning to demonstrate the defense effectiveness, as FUIA against sample and client unlearning both rely on GIA.